OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)?

OpenStudy (benlindquist):

get a line chart or whatever u need

OpenStudy (benlindquist):

liike a graph

pooja195 (pooja195):

Hi welcome to openstudy :) Use the slope formula to figure out the slope here it is : \[\LARGE~ \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE~ \frac{1--5}{4-1}=slope\] \[\LARGE~ \frac{1+5}{4-1}=slope\]